OpenStudy (anonymous):

Use three trapezoids of equal width to approximate the integral of lnx between 1 and 4. (1/2) + ln 12 ln 24 ln 12 (1/2) ln 24

OpenStudy (anonymous):

\[\int_1^4\ln x~dx\approx\frac{1}{2}(0+\ln 2)+\frac{1}{2}(\ln 2+\ln 3)+\frac{1}{2}(\ln3+\ln4)\] Simplify the right side and you're done. (The height of each trapezoid is implied, since it's just 1.)

OpenStudy (anonymous):

ln12?

OpenStudy (anonymous):

\[\frac{1}{2}\ln 2+\frac{1}{2}\ln 2+\frac{1}{2}\ln 3+\frac{1}{2}\ln 3+\frac{1}{2}\ln 4\\ \ln 2+\ln 3+\frac{1}{2}\ln 2^2\\ \ln 2+\ln 3+\ln 2\\ 2\ln 2+\ln 3\\ \ln 2^2+\ln 3\\ \ln 4+\ln 3\\ \ln12\] Yup, that's right